OpenStudy (anonymous):

verification please. (posted this b4 apologize)

OpenStudy (anonymous):

A rope is 40 meters long and (the rope) has the mass of 10kg. It hangs from one end off the tall building (the building is more than 40 meters). How much work is required to pull the rope up to the top of the building? Solution: Force = Mass * Acceleration due to gravity F = 10kg * 9.8m/s^2 = 98 Newtons. The distance is varying from y=0 to y=40, so our distance is\[\int\limits_{0}^{40}x~dx\] then Work = Force * Distance Work = \[98 \times \int\limits_{0}^{40}x~dx\]

OpenStudy (anonymous):

then I get:\[49x^2|^{40}_{0}~=~49 \times (40^2)~=~78,400\]

OpenStudy (anonymous):

so my answer is 78,400 joules.

OpenStudy (anonymous):

I said the distance is from y=0 to y=40 and then integrated with respect to x. My error... but am I getting the point ?

OpenStudy (anonymous):

why are you dividing by 40 ?

OpenStudy (perl):

we can assume the rope has uniform density

OpenStudy (perl):

the rope is 10 kg, and has length 40 meters density is 10 kg / 40 meter = 0.25 kg / meter

OpenStudy (perl):

linear density* i should say

OpenStudy (perl):

it also works out using center of mass approach. the center of mass is at height 20 meters, it has mass 10 kg, and you are moving the center of mass 20 meters. work = m* g * h = 10 * 9.8 * 20 = 1960 Joules

OpenStudy (perl):

check $$ \Large \int_{0}^{40}\frac{10}{40}y \cdot 9.8 dy = 1960 $$

OpenStudy (anonymous):

I don't know anything about the density, my teacher hasn't mentioned anything about it in class ( -:( ) I was just thinking that W = D * F And D= integral of x from 0 to 40, F= Mass times acceleration = 10 * 9.8 \[W(x)=(9.8 \times 10)\int\limits_{0}^{40}x~dx\] in class, we didn't talk much about density. Also, since the acceleration was given in the question, then be there density it would be given to me as well...

OpenStudy (anonymous):

the fact that density is not given to me, I guess, means that I don't need to worry about it.

OpenStudy (anonymous):

I am not trying to challenge you, I am just looking for understanding.

OpenStudy (anonymous):

you said, the rope is 10 kg, and has length 40 meters density is 10 kg / 40 meter But how does this "density" work out with that F = m * a W = F * D ?

OpenStudy (perl):

delta W = 10 kg / (40 meter) * (delta y meter) * 9.8 meter / second ^2 * y meter Joule has units kg * m^2 /s^2

OpenStudy (anonymous):

density * acceleration * distance, basically ?

OpenStudy (perl):

you can divide the rope into small chunks of width delta y

OpenStudy (anonymous):

yes, that is the only part that I get 100%. (Distance is an integral, because it varies) Kind of struggling on the other parts.

OpenStudy (anonymous):

cup ?

OpenStudy (anonymous):

well

OpenStudy (perl):



OpenStudy (anonymous):